# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v2

@@ Abstract
AI models make mistakes when recognizing images—whether in-domain, out-of-domain, or adversarial. Predicting these errors is critical for improving system reliability, reducing costly mistakes, and enabling proactive corrections in real-world applications such as healthcare, finance, and autonomous systems. However, understanding what mistakes AI models make, why they occur, and how to predict them remains an open challenge. Here, we conduct comprehensive empirical evaluations using a “mentor” model—a deep neural network designed to predict another “mentee” model’s errors. Our findings show that the mentor excels at learning from a mentee’s mistakes on adversarial images with small perturbations and generalizes effectively to predict in-domain and out-of-domain errors of the mentee. Additionally, transformer-based mentor models excel at predicting errors across various mentee architectures. Subsequently, we draw insights from these observations and develop an “oracle” mentor model, dubbed SuperMentor, that can outperform baseline mentors in predicting errors across different error types from the ImageNet-1K dataset. Our framework paves the way for future research on anticipating and correcting AI model behaviors, ultimately increasing trust in AI systems. All code, models, and data will be made publicly available.
@@ 1 Introduction
Subsequent research in out-of-domain detection has demonstrated that a model’s accuracy is often correlated with how far the data deviates from in-domain samples (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017). These methods typically rely on predefined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004) and confidence scores (Guillory et al., 2021), which limits their ability to generalize predictions across various data types, including errors arising from in-domain data or adversarial attacks (Szegedy, 2013). In parallel, earlier efforts in trustworthiness prediction mostly depend on shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018) or carefully-designed loss functions (Luo et al., 2021). However, they do not investigate how different error types influence trustworthiness prediction.
@@ 2 Related work
With the growing deployment of AI models across diverse fields, ensuring their reliability and understanding their limitations has become increasingly crucial. This has led to numerous research in safe AI such as uncertainty estimation (Nado et al., 2021; Lakshminarayanan et al., 2017), anomaly detection (Bogdoll et al., 2022), outlier detection (Boukerche et al., 2020) and out-of-domain detection (Yang et al., 2024). Unlike these areas, which mainly aim to predict whether the input data falls outside the training domain, our focus is on monitoring and predicting errors in AI models by determining whether the model’s output is correct, irrespective of whether the data comes from the training domain. To detect whether the input data is out of scope, the prior approaches mainly rely on softmax outputs (Granese et al., 2021; Hendrycks & Gimpel, 2016; Dang et al., 2024), activations from network layers (Wang et al., 2020; Cheng et al., 2019; Ferreira et al., 2023), shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018), and carefully-designed loss functions (Luo et al., 2021) in applications such as object detection (Kang et al., 2018) and trajectory prediction (Shao et al., 2023, 2024). However, these methods often rely on manually defined metrics to estimate the likelihood of a mentee making mistakes, or they fail to examine how different error types affect error prediction. In contrast, our strategy employs a separate deep neural network to automatically learn and approximate the mentee’s decision boundaries for specific error types, providing an end-to-end trainable framework for error prediction. Moreover, our research direction differs from selective prediction (Geifman & El-Yaniv, 2017), rejection learning (Cortes et al., 2016), and learning to defer (Madras et al., 2018) by avoiding joint training with the mentee and requiring no knowledge of the mentee’s training data. Besides, although rejecting a mentee’s unreliable predictions and predicting the correctness of a mentee’s outputs are often positively correlated, they remain distinct tasks.
@@
Recent studies (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017) have shown that a model’s accuracy on a given dataset is often correlated with how far the data deviates from in-domain samples. However, these studies typically rely on pre-defined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004), confidence scores (Guillory et al., 2021), domain-invariant representations (Chuang et al., 2020), and domain augmentation (Deng et al., 2021a), limiting their ability to generalize error prediction beyond in-domain data. In contrast, our mentor is capable of predicting both OOD and in-domain errors for a mentee. Additionally, our mentor is an AI model trained end-to-end without relying on manually defined criteria.
@@
In addition to OOD error, (Szegedy, 2013) discovered that deep neural networks can be fooled using input perturbations of extremely low magnitude. Building upon this finding, a substantial number of adversarial attacks have been proposed, including white-box attacks (Goodfellow et al., 2014; Mądry et al., 2017; Carlini & Wagner, 2017; Schwinn et al., 2023; Gao et al., 2020), black-box attacks (Uesato et al., 2018; Rahmati et al., 2020; Brendel et al., 2017; Chen et al., 2020), and backdoor attacks (Liu et al., 2020b; Xie et al., 2019; Kolouri et al., 2020). To defend against these adversarial attacks, various defence mechanisms (Qin et al., 2019; Deng et al., 2021b; Liu et al., 2019) have been developed to withstand or detect adversarial inputs. Furthermore, although the primary objective of adversarial attacks is to deceive AI models, there are instances where adversarial perturbations are exploited to enhance the model performance — a technique known as adversarial training (Ilyas et al., 2019; Gowal et al., 2020; Balunović & Vechev, 2020). Unlike adversarial training, which involves using adversarial samples to train the mentee, our approach focuses on teaching mentors to learn the mentee’s error patterns revealed by these adversarial attack samples.
@@ 3.1 Mentors
Model architecture: We propose mentor models, as illustrated in Fig. 2. Given an input image, the backbone of a mentor model extracts features from the input image. We adopt either of the two backbones for the feature extractors of mentors: a 2D Convolutional Neural Network (2D-CNN) ResNet50 (He et al., 2016) and a transformer-based ViT (Dosovitskiy, 2020). The extracted feature maps are further processed in two streams implemented as multi-layer perceptrons (MLP)s. The parameters of the MLPs in the two streams are not shared.
The first stream generates logits by predicting the probability distribution of a mentee over all the object classes when the mentee classifies the given image. The mentee network is kept fixed while training the mentor. Let us define the mentee’s output logit as . We introduce the distillation loss proposed by (Hinton, 2015): to align with , where represents Kullback-Leibler divergence. is the softmax function. is the temperature, which controls the smoothness of the soft probability distribution.
In the second stream, the mentor is prompted to predict whether the mentee will make a mistake on the given image or not. We denote the predicted binary label as , where 1 indicates that the mentee does not make a mistake and vice versa for 0. This prediction is supervised by , where is the ground truth correctness label of a mentee and represents the mentor’s predicted probability that the mentee’s prediction is correct. The overall loss is .
Training and implementation details: All mentors are trained on Nvidia RTX A5000 and A6000 GPUs, utilizing the AdamW optimizer (Loshchilov & Hutter, 2017) with a cosine annealing scheduler (Loshchilov & Hutter, 2016), and an initial learning rate of . All mentors load the weights of the feature extractor pre-trained on the ImageNet-1K dataset for 1000-way image classification tasks (Deng et al., 2009) and further fine-tune on the error prediction task. During training, images are resized and center-cropped to pixels. All the mentor models are trained for 40 epochs with a batch size of 512.
@@ 3.2 Mentees and their datasets
We employ two architectures as the mentees’ backbones: ResNet50 (He et al., 2016), which is a 2D-Convolutional neural network (2D-CNN), and ViT (Dosovitskiy, 2020), which is a transformer based on self-attention mechanisms.
@@ 3.3 Datasets for training and testing mentors
In-Domain (ID) Errors occur on data that come from the same domain as the mentee’s training dataset. Specifically, errors on images from the standard validation set of ImageNet-1K or the test sets of CIFAR-10 and CIFAR-100 are considered ID errors. Along with the correctly classified images from these standard test sets, we create three datasets for a mentor: IN-ID, C10-ID, and C100-ID, following the naming convention of [Dataset]-[Error Type].
@@
Training and test splits. The number of correctly and incorrectly classified images by a mentee can vary significantly based on its classification performance. To create balanced training and test sets for the mentor and avoid the long-tailed distribution problem, we select an equal number of correctly and incorrectly classified samples for each training batch and every test set. See Appendix, Sec. A for more details.
@@ 3.4 Baselines and evaluation metric
Baselines. We include seven baseline methods for error prediction of a mentee. 1) Self Error Rate (SER) predicts the correctness of the mentee’s outputs by randomly assigning “correct” or “wrong” based on the mentee’s in-domain accuracy; 2) Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2016) classify the mentee’s predictions as correct if its MCP exceeds a predefined threshold. 3) Class Probability Entropy (CPE) evaluates a mentee’s prediction as correct if its entropy of the probability distribution over all classes is below a predefined threshold. 4) Distance To Centroid (DTC) considers a prediction correct if its feature embedding is within a predefined distance from the class centroid. 5) ConfidNet (Corbière et al., 2019) predicts failure by learning True Class Probability (TCP) on the training samples. 6) TrustScore (Jiang et al., 2018) employs Trust Score to quantify the agreement between the classifier and a modified nearest-neighbor classifier. 7) Steep Slope Loss (SSL) (Luo et al., 2021) develops a neural network for predicting trustworthiness by utilizing a steep slope loss function. See Appendix, Sec. B for more details.
Evaluation metric. To assess the performance of mentors, we report their error prediction accuracy on the test set corresponding to each specified error source. For instance, a mentor trained on the C10-ID training set is evaluated on the C10-OOD-SpN test set. The error prediction accuracy is calculated by averaging the mentor’s accuracies on the samples that the mentee correctly classified and those that the mentee incorrectly classified. However, since a mentee can make mistakes across various real-world scenarios, a mentor must accurately predict errors across all error types. Therefore, we compute the average accuracy, named as Accuracy (%) of a mentor across all test sets, including one ID error, four OOD errors, and four AA errors. See Appendix, Sec. C for an example calculation of the average accuracy of a mentor.
@@ 4.1 Training on specific errors of mentees impacts the performance of mentors
A mentee’s mistakes can reveal their learning biases, behaviors, or traits. Here, we investigate which types of errors offer the most insight into understanding a mentee’s decision boundaries during image recognition tasks. We train mentors with identical architectures on datasets containing specific error types made by the mentee across C10 (Fig. 3(a)), C100 (Fig. 3(b)), and IN (Fig. 3(c)). For instance, if a mentor trained on C10-OOD achieves higher accuracy in error prediction compared to one trained on C10-IN, this suggests that in-domain errors provide less diagnostic information about the mentee’s decision-making process than out-of-domain errors. Both mentors and mentees may have the same or different backbones, such as ResNet50 (R) or ViT (V).
As shown in Fig. 3, over C10, C100, and IN images, the high accuracy for mentors trained on adversarial attack (AA) errors indicates that these errors offer deeper insights into the mentee’s decision process compared to out-of-domain (OOD) and in-domain (ID) errors. In some cases, mentors trained on OOD errors slightly outperformed those trained on ID errors, though both were still inferior to those trained on AA errors. Interestingly, our experiments reveal that training mentors exclusively on AA errors performs marginally worse than joint training on ID, OOD, and AA errors (see Appendix, Sec. D). This result underscores a significant overlap of mentee error patterns among ID, OOD, and AA data.
Loss landscape analysis. A loss landscape of a mentee reflects how a mentee’s loss function behaves across different parameter configurations. Mentors’ performance offers insights into the structure of a mentee’s loss landscape. Consistent with (Ilyas et al., 2019), the high accuracy of mentors trained on AA errors suggests that adversarial images lie closer to the mentee’s decision boundary, enabling more accurate prediction of the mentee’s mistakes and a deeper understanding of the loss landscape. Similarly, OOD data aids mentors in learning boundaries by shifting ID samples closer to the boundary. However, it does not explore the boundary as thoroughly as adversarial images. ID data, with fewer samples near the boundary, provides more limited exploration compared to adversarial images.

### v3

@@ Abstract
AI models make mistakes when recognizing images—whether in-domain, out-of-domain, or adversarial. Predicting these errors is critical for improving system reliability, reducing costly mistakes, and enabling proactive corrections in real-world applications such as healthcare, finance, and autonomous systems. However, understanding what mistakes AI models make, why they occur, and how to predict them remains an open challenge. Here, we conduct comprehensive empirical evaluations using a “mentor” model—a deep neural network designed to predict another “mentee” model’s errors. Our findings show that the mentor excels at learning from a mentee’s mistakes on adversarial images with small perturbations and generalizes effectively to predict in-domain and out-of-domain errors of the mentee. Additionally, transformer-based mentor models excel at predicting errors across various mentee architectures. Subsequently, we draw insights from these observations and develop an “oracle” mentor model, dubbed SuperMentor, that can outperform baseline mentors in predicting errors across different error types from the ImageNet-1K dataset. Our framework paves the way for future research on anticipating and correcting AI model behaviors, ultimately increasing trust in AI systems. Our data and code are available at here.
@@ 1 Introduction
Subsequent research in out-of-domain detection has demonstrated that a model’s accuracy is often correlated with how far the data deviates from in-domain samples (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017). These methods typically rely on predefined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2022; Madani et al., 2004) and confidence scores (Guillory et al., 2021), which limits their ability to generalize predictions across various data types, including errors arising from in-domain data or adversarial attacks (Szegedy et al., 2014). In parallel, earlier efforts in trustworthiness prediction mostly depend on shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018) or carefully-designed loss functions (Luo et al., 2021). However, they do not investigate how different error types influence trustworthiness prediction.
@@ 2 Related work
With the growing deployment of AI models across diverse fields, ensuring their reliability and understanding their limitations has become increasingly crucial. This has led to numerous research in safe AI such as uncertainty estimation (Nado et al., 2021; Lakshminarayanan et al., 2017), anomaly detection (Bogdoll et al., 2022), outlier detection (Boukerche et al., 2020) and out-of-domain detection (Yang et al., 2024). Unlike these areas, which mainly aim to predict whether the input data falls outside the training domain, our focus is on monitoring and predicting errors in AI models by determining whether the model’s output is correct, irrespective of whether the data comes from the training domain. To detect whether the input data is out of scope, the prior approaches mainly rely on softmax outputs (Granese et al., 2021; Hendrycks & Gimpel, 2017; DANG et al., ), activations from network layers (Wang et al., 2020; Cheng et al., 2019; Ferreira et al., 2023), shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018), and carefully-designed loss functions (Luo et al., 2021) in applications such as object detection (Kang et al., 2018) and trajectory prediction (Shao et al., 2023, 2024). However, these methods often rely on manually defined metrics to estimate the likelihood of a mentee making mistakes, or they fail to examine how different error types affect error prediction. In contrast, our strategy employs a separate deep neural network to automatically learn and approximate the mentee’s decision boundaries for specific error types, providing an end-to-end trainable framework for error prediction. Moreover, our research direction differs from selective prediction (Geifman & El-Yaniv, 2017), rejection learning (Cortes et al., 2016), and learning to defer (Madras et al., 2018) by avoiding joint training with the mentee and requiring no knowledge of the mentee’s training data. Besides, although rejecting a mentee’s unreliable predictions and predicting the correctness of a mentee’s outputs are often positively correlated, they remain distinct tasks.
@@
Recent studies (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017) have shown that a model’s accuracy on a given dataset is often correlated with how far the data deviates from in-domain samples. However, these studies typically rely on pre-defined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2022; Madani et al., 2004), confidence scores (Guillory et al., 2021), domain-invariant representations (Chuang et al., 2020), and domain augmentation (Deng et al., 2021a), limiting their ability to generalize error prediction beyond in-domain data. In contrast, our mentor is capable of predicting both OOD and in-domain errors for a mentee. Additionally, our mentor is an AI model trained end-to-end without relying on manually defined criteria.
@@
In addition to OOD error, (Szegedy et al., 2014) discovered that deep neural networks can be fooled using input perturbations of extremely low magnitude. Building upon this finding, a substantial number of adversarial attacks have been proposed, including white-box attacks (Goodfellow et al., 2015; Mądry et al., 2017; Carlini & Wagner, 2017; Schwinn et al., 2023; Gao et al., 2020), black-box attacks (Uesato et al., 2018; Rahmati et al., 2020; Brendel et al., 2021; Chen et al., 2020), and backdoor attacks (Liu et al., 2020b; Xie et al., 2019; Kolouri et al., 2020). To defend against these adversarial attacks, various defence mechanisms (Qin et al., 2020; Deng et al., 2021b; Liu et al., 2019) have been developed to withstand or detect adversarial inputs. Furthermore, although the primary objective of adversarial attacks is to deceive AI models, there are instances where adversarial perturbations are exploited to enhance the model performance — a technique known as adversarial training (Ilyas et al., 2019; Gowal et al., 2020; Balunović & Vechev, 2020). Unlike adversarial training, which involves using adversarial samples to train the mentee, our approach focuses on teaching mentors to learn the mentee’s error patterns revealed by these adversarial attack samples.
@@ 3.1 Mentors
Model architecture: We propose mentor models, as illustrated in Fig. 2. Given an input image, the backbone of a mentor model extracts features from the input image. We adopt either of the two backbones for the feature extractors of mentors: a 2D Convolutional Neural Network (2D-CNN) ResNet50 (He et al., 2016) and a transformer-based ViT (Dosovitskiy et al., 2020). The extracted feature maps are further processed in two streams implemented as multi-layer perceptrons (MLP)s. The parameters of the MLPs in the two streams are not shared.
The first stream generates logits by predicting the probability distribution of a mentee over all the object classes when the mentee classifies the given image. The mentee network is kept fixed while training the mentor. Let us define the mentee’s output logit as . We introduce the distillation loss proposed by (Hinton et al., 2015): to align with , where represents Kullback-Leibler divergence. is the softmax function. is the temperature, which controls the smoothness of the soft probability distribution.
In the second stream, the mentor is prompted to predict whether the mentee will make a mistake on the given image or not. We denote the predicted binary label as , where 1 indicates that the mentee does not make a mistake and vice versa for 0. This prediction is supervised by , where is the ground truth correctness label of a mentee and represents the mentor’s predicted probability that the mentee’s prediction is correct. The overall loss is , where controls the dynamic weighting between and over training epochs. Here, denotes the current training epoch, is the total number of training epochs, and governs the rate at which evolves throughout training.
Training and implementation details: All mentors are trained on NVIDIA RTX A6000 GPUs, utilizing the AdamW optimizer (Loshchilov & Hutter, 2019) with a cosine annealing scheduler (Loshchilov & Hutter, 2022). The initial learning rate is set to for ResNet50 mentors and for ViT mentors. All mentors load the weights of the feature extractor pre-trained on the ImageNet-1K dataset for 1000-way image classification tasks (Deng et al., 2009) and further fine-tune on the error prediction task. During training, images are resized and center-cropped to pixels. All mentors are trained for 30 epochs with a batch size of 32 on the CIFAR-10 and CIFAR-100 and 384 on the ImageNet-1K. The value of in the loss function is chosen for mentors to approach optimal performance. Specifically, is set to 1.0 for the mentors on CIFAR-10, 2.0 for those on CIFAR-100, 3.0 for ResNet50 mentors on ImageNet-1K (Deng et al., 2009), and 0.1 for ViT mentors on ImageNet-1K.
@@ 3.2 Mentees and their datasets
We employ two architectures as the mentees’ backbones: ResNet50 (He et al., 2016), which is a 2D-Convolutional neural network (2D-CNN), and ViT (Dosovitskiy et al., 2020), which is a transformer based on self-attention mechanisms.
@@ 3.3 Datasets for training and testing mentors
In-Domain (ID) Errors occur on data that come from the same domain as the mentee’s training dataset. Specifically, errors on images from the standard validation set of IN or the test sets of C10 and C100 are considered ID errors. Along with the correctly classified images from these standard test sets, we create three datasets for a mentor: IN-ID, C10-ID, and C100-ID, following the naming convention of [Dataset]-[Error Type].
@@
Training and test splits for mentors. Given the three types of errors described above, we partition the data for mentors into training and testing sets using a 70/30 split. For example, 70% of the original C10 evaluation images, along with their corresponding error-modified versions based on mentee performance, are used for training, while the remaining 30% are reserved for testing. This ensures that mentor models are trained and evaluated on samples derived from the same set of original images, without any overlap in the original image content, but with distinct domain shifts. The same strategy is consistently applied to the C100 and IN datasets. To mitigate the effects of long-tailed class distributions, each training batch includes an equal number of correctly and incorrectly classified samples.
@@ 3.4 Baselines and evaluation metric
Baselines. We include seven baseline methods for error prediction of a mentee. 1) Self Error Rate (SER) predicts the correctness of the mentee’s outputs by randomly assigning “correct” or “wrong” based on the mentee’s in-domain accuracy; 2) Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2017) classify the mentee’s predictions as correct if its MCP exceeds a predefined threshold. 3) Class Probability Entropy (CPE) evaluates a mentee’s prediction as correct if its entropy of the probability distribution over all classes is below a predefined threshold. 4) Distance To Centroid (DTC) considers a prediction correct if its feature embedding is within a predefined distance from the class centroid. 5) ConfidNet (Corbière et al., 2019) predicts failure by learning True Class Probability (TCP) on the training samples. 6) TrustScore (Jiang et al., 2018) employs Trust Score to quantify the agreement between the classifier and a modified nearest-neighbor classifier. 7) Steep Slope Loss (SSL) (Luo et al., 2021) develops a neural network for predicting trustworthiness by utilizing a steep slope loss function. See Appendix, Sec. A for more details.
Evaluation metric. To assess the performance of mentors, we report their error prediction accuracy on the test set corresponding to each specified error source. For instance, a mentor trained on the C10-ID training set is evaluated on the C10-OOD-SpN test set. The error prediction accuracy is calculated by averaging the mentor’s accuracies on the samples that the mentee correctly classified and those that the mentee incorrectly classified. However, since a mentee can make mistakes across various real-world scenarios, a mentor must accurately predict errors across all error types. Therefore, we compute the average accuracy, named as Accuracy (%) of a mentor across all test sets, including one ID error, four OOD errors, and four AA errors. See Appendix, Sec. B for an example calculation of the average accuracy of a mentor.
@@ 4.1 Training on specific errors of mentees impacts the performance of mentors
A mentee’s mistakes can reveal their learning biases, behaviors, or traits. Here, we investigate which types of errors offer the most insight into understanding a mentee’s decision boundaries during image recognition tasks. We train mentors with identical architectures on datasets containing specific error types made by the mentee across C10 (Fig. 3(a)), C100 (Fig. 3(b)), and IN (Fig. 3(c)). For instance, if a mentor trained on C10-OOD achieves higher accuracy in error prediction compared to one trained on C10-ID, this suggests that in-domain errors provide less diagnostic information about the mentee’s decision-making process than out-of-domain errors. Both mentors and mentees may have the same or different backbones, such as ResNet50 (R) or ViT (V).
As shown in Fig. 3, over C10, C100, and IN images, the high average accuracies for mentors trained on adversarial attack (AA) errors indicates that these AA errors offer deeper insights into the mentee’s decision process compared to out-of-domain (OOD) and in-domain (ID) errors. To validate this, we first conducted an ANOVA test on the three groups of results (ID, OOD, and AA) for each of the three datasets (C10, C100, and IN). In all cases, the p-values are below 0.05, indicating a statistically significant difference in error prediction performance among mentors trained on different error types. Next, we performed two-tailed t-tests to compare AA-trained mentors against ID-trained and OOD-trained mentors across four [mentee]-[mentor] configurations: R–R, V–R, R–V, and V–V. All pairwise p-values are below 0.05, confirming that AA-trained mentors significantly outperform those trained on ID or OOD error types. In addition, it is worth noting that, in some cases, mentors trained on OOD errors slightly outperformed those trained on ID errors, although both still performed worse than those trained on AA errors. Interestingly, our experiments reveal that training mentors exclusively on AA errors performs marginally worse than joint training on all three error types (see Appendix, Sec. C). This result underscores a significant overlap of mentee error patterns among ID, OOD, and AA data.
Loss landscape analysis. A loss landscape of a mentee reflects how a mentee’s loss function behaves across different parameter configurations. Mentors’ performance offers insights into the structure of a mentee’s loss landscape. Consistent with (Ilyas et al., 2019), the high accuracy of mentors trained on AA errors suggests that adversarial images lie closer to the mentee’s decision boundary, enabling more accurate prediction of the mentee’s mistakes and a deeper understanding of the loss landscape. Similarly, OOD data aids mentors in learning boundaries by shifting ID samples closer to the boundary. However, it does not explore the boundary as thoroughly as adversarial images. ID data, with fewer samples near the boundary, provides more limited exploration compared to adversarial images. To support our claims, we include quantitative loss landscape analyses in Appendix, Sec. D. The results show that mentors trained on AA errors exhibit a wider loss landscape compared to those trained on ID or OOD errors.